OpenStudy (lilkg77):

The first quartile of a data set is 72, and the third quartile is 92. Which of these values in the data set is an outlier? A. 121 B. 101 C. 41 D. 61

OpenStudy (ageta):

The interquartile range is used to check for the outliers present in the data set. A data value x is considered as an outlier, if x < Q1 - 1.5(IQR) or x > Q3 + 1.5(IQR). this is what i found in my notes but idk how to calculate it one sec

OpenStudy (ageta):

can u plug thst in this eqn x < Q1 - 1.5(IQR) or x > Q3 + 1.5(IQR). to find outliers?

OpenStudy (ageta):

you know what the IQR is right?

OpenStudy (lilkg77):

Quartile range right?

OpenStudy (ageta):

yes

OpenStudy (ageta):

i have to go

OpenStudy (ageta):

bumb your question to the top so other can help you bye

OpenStudy (anonymous):

i here

OpenStudy (lilkg77):

Show me how clever u are @TheCleverOne

OpenStudy (anonymous):

how

OpenStudy (anonymous):

lilkg77 its me milk333

OpenStudy (anonymous):

lilkg77 its me milk333

OpenStudy (lilkg77):

Oh whats up @TheCleverOne

OpenStudy (lilkg77):

Changed users?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

because someone else had the name TheSmartOne and it annoyed me so i changed and thesmartone is not smart

OpenStudy (lilkg77):

Oh lol @TheCleverOne

OpenStudy (lilkg77):

Think u could help with this question?

OpenStudy (anonymous):

lets see...

jimthompson5910 (jim_thompson5910):

use the formulas @ageta wrote at the top of the page

jimthompson5910 (jim_thompson5910):

IQR = Q3 - Q1 = 92 - 72 = 20 what is `Q1 - 1.5(IQR)` equal to?